### v3

# An open dataset for oracle bone script recognition and decipherment

Abstract
Oracle Bone Script (OBS), one of the earliest known forms of ancient Chinese writing, holds invaluable insights into the humanities and geography of the Shang Dynasty, dating back 3,000 years. The immense historical and cultural significance of these writings cannot be overstated. However, the passage of time has obscured much of their meaning, presenting a significant challenge in deciphering these ancient texts. With the advent of Artificial Intelligence (AI), employing AI to assist in interpreting OBS has become a feasible option. Yet, progress in this area has been hindered by a lack of high-quality datasets. To address this issue, this paper details the creation of the HUST-OBC dataset. This dataset encompasses 77,064 images of 1,588 individual deciphered scripts and 62,989 images of 9,411 undeciphered characters, with a total of 140,053 images, compiled from diverse sources. Additionally, all images and labels have been reviewed and corrected by experts in oracle bone studies. The hope is that this dataset could inspire and assist future research in deciphering those unknown OBS. All the codes and datasets are available at https://github.com/Pengjie-W/HUST-OBC.
Background & Summary
Figure 1: Archaeological discovery of 3000-year-old Shang Dynasty oracle bones.
Oracle Bone Script (OBS), etched onto turtle shells and animal bones, stands as one of the earliest forms of writing discovered in China (see Figure 1). These inscriptions, dating back about 3,000 years, offer a window into the human geography of the Shang Dynasty, an ancient feudal era. The content encompasses a range of topics including astrology, meteorology, animal husbandry, religion, and ritual practices [1, 2]. Similar to other ancient scripts, the meanings of many OBS have been lost over time. Of the 160,000 pieces unearthed, they reveal more than 4,600 distinct oracle bone characters, yet only about a thousand of these have been deciphered, with their meanings and corresponding modern Chinese characters confirmed [3].
The task of deciphering these ancient scripts is complicated by several factors. Historically, the methods of preservation and excavation were not always ideal, leading to many of the oracle bones being damaged. This damage often results in partial, unclear, or illegible inscriptions, making interpreting them arduous. In addition, the nature of OBS as an early writing system means that it underwent significant evolution. There is a considerable variation in the form of characters, with many characters appearing in multiple, sometimes radically different forms [4]. This variability adds another layer of complexity to the deciphering process. All these factors contribute to making the full understanding of OBS not only challenging but also a rare feat, attracting the keen interest of scholars and historians alike in the field of ancient Chinese studies.
Over the past decades, it has witnessed the widespread applications of Artificial Intelligence (AI) in various fields. Notably, the immense success of Optical Character Recognition (OCR) technology in processing modern texts has sparked interest in the potential use of AI to aid in deciphering OBS. Modern AI algorithms, particularly those centered around deep learning models like artificial neural networks, typically require an extensive volume of data for training. This approach enables them to achieve, and sometimes surpass, human-level performance in specific tasks. A fundamental step towards employing these models to decipher OBS involves the creation and annotation of a comprehensive high-quality dataset of OBS. There have been some pioneering efforts in this area. For instance, Li et al. [5] built the HWOBC dataset by engaging experts from diverse academic backgrounds to manually write OBS and scan the images. Fu et al. [6] and Yue et al. [7] subsequently proposed the OBI-100 and OBI-125 datasets, with the images scanned from books related to OBS research. Additionally, Guo et al. [8] collected more than 20k OBS images from various websites to build the Oracle-20k dataset. These efforts lay a solid foundation for digitalization and recognition research of OBS. However, the fact that these datasets have certain limitations also hinders their use in AI-assisted OBS decipherment:
  • They often have limited categories and samples of OBS due to data collection from a single source.
  • The annotation of the categories might not be deduplicated, resulting in the same OBS being categorized into different classes.
  • The lack of cross-validation from multiple sources casts doubt on the accuracy of some data.
  • The datasets comprise only deciphered oracle bone images, making them unsuitable for deciphering tasks.
  • Some datasets contain unprocessed images, filled with noise or blur.
To address these issues, we propose the high-quality HUST-OBC dataset. The HUST-OBC dataset was collected from three different sources, including books, websites, and existing databases. It includes images of OBS transcribed by experts as well as images cropped from actual photos of oracle bones, ensuring diversity. To build HUST-OBC, we designed a semi-automatic pipeline that collects and annotates data from various sources and had OBS experts review and calibrate the final dataset. As shown in Table 1, HUST-OBC contains over 10k deciphered and undeciphered OBS categories and more than 140k images, making it one of the largest datasets for OBS recognition and deciphering to date. We hope HUST-OBC will aid and inspire future AI-assisted OBS research.
Dataset Image Source Deciphered Undeciphered
#Characters #Images #Characters #Images
HWOBC Book 1,566 33,555 2,315 49,688
OBI-100 Book 100 4,748 0 0
OBI-125 Book 360 4,725 0 0
Oracle-20k Web 261 20,039 0 0
HUST-OBC Book, Web, Database 1,588 77,064 9,411 62,989
Table 1: Comparison of the scale of HUST-OBC with other oracle bone script datasets.
Methods
To construct a diverse dataset, we gathered images of OBS from three distinct sources: book, website, and database. To organize and merge data from these varied origins, as shown in Figure 2, we designed a semi-automated pipeline comprising four key steps: Data Acquisition, Automatic Annotation, Data Integration, and Data Validation. In this section, we will delve into the details of each step.
Data Acquisition
OBS is inscribed on turtle shells and animal bones and boasts a history of over 3,000 years. These precious artifacts are dispersed in museums and private collections worldwide, where they are meticulously preserved, making direct access to the text inscribed on the original oracle bones quite challenging. Thankfully, most of the publicly available oracle bones have been replicated or transcribed by experts, making them accessible in various forms for scholarly research. As illustrated in Table 3, the HUST-OBC dataset is constructed by gathering data from these diverse sources.
Figure 2: Flowchart of building the HUST-OBC dataset.
Books
Books remain the predominant form in compiling oracle bone scripts, with most discovered scripts to date collected and interpreted in volumes like the New Compilation of Oracle Bone Scripts, which ensures accuracy by incorporating the latest research in this field. Specifically, we utilized the following books as data sources while constructing the HUST-OBC dataset.
A. New Compilation of Oracle Bone Scripts ({CJK*}UTF8gbsn新甲骨文编) aaahttps://books.google.com/books?id=S0RergEACAAJ encompasses samples of Oracle Bone Script found since its initial discovery, as presented in all public materials.
B. Oracle Bone Script: Six Digit Numerical Code ({CJK*}UTF8gbsn甲骨文六位数字码检索字库) bbbhttps://www.amazon.com/Oracle-digit-search-character-Chinese/dp/7557901509 assigns digital codes to oracle bone scripts, annotating each code with its corresponding oracle bone character, modern Chinese character form, provenance, and other relevant details.
Since these books do not provide electronic databases or original image data, we manually scanned the pages of these books, obtaining 1,054 and 700 pages from books A and B, respectively. An example of scanned pages is presented in Figure 3(a).
Websites
With the widespread adoption of the Internet, websites have emerged as an alternative for hosting oracle bone data, offering more convenient retrieval capabilities. We have designed a web crawler program to collect data from the following websites:
C. GuoXueDaShi ({CJK*}UTF8gbsn国学大师) ccchttps://www.guoxuedashi.net/jgwhj/, initiated and maintained by enthusiasts of Chinese classical studies, which includes various historical texts including dictionaries, histories, etc. A screenshot of the website is shown in Figure 4(b).
D. YinQiWenYuan ({CJK*}UTF8gbsn殷契文渊) dddhttps://jgw.aynu.edu.cn is a data platform maintained by the Key Laboratory of Oracle Bone Inscriptions Information Processing, archives various types of data, including photos of the original oracle bones, transcribed characters, and related research articles. A screenshot of the website is shown in Figure 4(a).
These websites feature well-organized collections of oracle bone script images, which have been meticulously scanned, cropped, and aligned. They are systematically categorized across various web pages, facilitating the use of web crawler technology to download these images in a categorized format efficiently.
Databases
In recent years, the digitalization of ancient manuscripts and advancements in handwritten text recognition technology have opened new avenues in the study of oracle bone script, which has led to the proposal of relevant datasets. We have included the following databases in HUST-OBC.
E. HWOBC eeehttps://jgw.aynu.edu.cn/DownPage is a database specifically designed for the study and recognition of OBS. It contains tens of thousands of images of oracle bone script, which are hand-written by scholars from various backgrounds, based on the photos of original oracle bone characters.
(a) Scanned Book Page
(b) Page Cropping
Figure 3: Extraction of oracle bone script from books.
(a) YinQiWenYuan
(b) GuoXueDaShi
Figure 4: Screenshots of example website pages from YinQiWenYuan and GuoXueDaShi.
Automatic Annotation
Through data acquisition, we have gathered raw data from diverse sources. However, this data, in its current assorted formats, is not immediately usable. Hence, it necessitates further processing, including tasks such as cropping, annotating, and filtering.
Books
The raw data for the books consists of scanned images of pages, each displaying several oracle bone scripts along with their corresponding annotations in modern standard Chinese. As shown in Figure 3(a), despite differences in the layout of the New Compilation of Oracle Bone Scripts (left) and Oracle Bone Script: Six Digit Numerical Code (right), they both employ a table-like vertical format. This arrangement facilitates the use of computer vision algorithms, such as edge detection, to automatically extract content from these pages. Specifically, as shown in Figure 3(b), we employed edge detection and other techniques from the OpenCV toolkit [9] to crop the original scanned images by oracle bone characters, thereby obtaining individual slices of these characters. These slices are then categorized according to the layout rules, with each assigned a corresponding category ID. For example, as illustrated on the left side of Figure 3(b), the top of each column in the book is marked with the modern Chinese character equivalent to the oracle bone script. If a column lacks such a marking, it implies that it belongs to the same category as the adjacent column on its right. In the figure, we used different colors of dashed lines to distinguish between categories. Using this method, we extracted 24,558 and 14,053 oracle bone script images from source A and B books, respectively.
Although the slices are grouped during the cropping process, the corresponding modern Chinese character of each category remains unknown. A straightforward solution to determine the specific characters for each category is to use OCR techniques to recognize the marks in the books. However, most of the off-the-shelf OCR engines were trained only on commonly used modern Chinese characters and struggled to recognize the LidingfffLiding ({CJK*}UTF8gbsn隶定) is the practice of rewriting ancient Chinese character forms in clerical or regular script. and unknown characters that may be presented in these books. To address this issue, we trained a category assigner (see Figure 5) to automatically identify these labels. The specific training procedures are detailed as follows:
Figure 5: Schematic of the proposed category assigner.
  1. 1.
    Training Data Generation: The Chinese character labels we need to identify are all in standard print typeface (as seen in the Chinese characters on the outside of the table at the top of Figure 3(b), left side), and each cropped image contains only a single, individual character. Thus, we can conveniently generate corresponding training samples using a similar SimSun font. As shown in the block on the left side of Figure 5, we generated font images for all realistic Chinese characters according to the Ideographic Description Sequence (IDS) and assigned each a unique category ID. Additionally, to address the recognition of uncommon characters that may appear in books, we randomly synthesized Liding text using components like radicals of Chinese characters to serve as the Liding category for training purposes. Practically, we generated one image for each of the total 88,899 Chinese characters included in the IDS and randomly synthesized $\alpha$ Liding character images in each training epoch.
  2. 2.
    Training: Since each sample image contains only a single character, it is sufficient to train a simple classifier for recognition. For this purpose, we employed ResNet-50 [10] as the backbone network to train the classifier. Additionally, we utilized a weighted balanced cross-entropy loss $L$ to address the issue of the imbalance in the number of training samples across different categories:
    $L=-\frac{1}{N+1}\left[\sum_{i=1}^{N}y_{i}\log\left(p_{i}\right)+\frac{1}{\alpha}y_{N+1}\log\left(p_{N+1}\right)\right]$ (1)
    where $N$ is the number of categories, $P_{i}$ and $Y_{i}$ respectively represent the probabilities of the predicted and true labels being the $\text{i}^{th}$ category, taking values of 1 or 0, and $\alpha$ is the number of synthesized Liding samples in each training epoch.
  3. 3.
    Inference: During the inference phase, we input the Chinese character label images, which are cropped from the books, into the classification model trained in the second step. This process helps us determine the corresponding Chinese character for each category ID or whether it is a Liding character.
After completing the aforementioned procedures, all oracle bone script images contained within the scanned pages of sourced books acquired during the data acquisition phase have been automatically extracted and accurately categorized according to their respective classes.
Modern Character Category Oracle Bone Script Images Liding Character Category Oracle Bone Script Images
[Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Table 2: Comparison of oracle bone script annotations using Modern and Liding characters as category labels.
Websites & Databases
The images of the oracle bone script collected from websites and databases have already been preprocessed by scanning, cropping, and alignment. Therefore, there is no need to design automatic annotation algorithms for this data, unlike the approach required for data from book sources. However, the following essential processes are still required:
  1. 1.
    Filtering: It is important to note that a significant portion of the data on the GuoXueDaShi website, contributed by enthusiasts of ancient Chinese culture, cannot be fully guaranteed for reliability. Consequently, in our filtration process, we cross-referenced these with other sources. This allowed us to identify 1,390 categories of oracle bone script images that were unique to GuoXueDaShi and could not be verified. As a result, we retained only 1,366 out of the initial 2,756 categories after excluding these unverifiable samples. The samples of these 1390 categories, due to their lack of reliability, have not been classified as deciphered or undeciphered samples and are stored separately in the dataset.
  2. 2.
    Code Matching: The oracle bone script images from online and database sources are marked with specific codes, which we further mapped into modern Chinese characters. For the oracle bone inscriptions of YinQiWenYuan and HWOBC, the HUST-OBC dataset only includes individual oracle bone characters, not compound characters. The term ’compound characters’ refers to oracle bone inscriptions corresponding to two or more words. Moreover, HWOBC is classified based on the character forms, leading to multiple character forms corresponding to the same Chinese character. Here, we merge them into the same category based on the corresponding Chinese character.
Figure 6: Oracle bone script images classified into the redundant category exhibit higher similarity in the feature vector space.
Integration
In the stages of Data Acquisition and Automatic Annotation, images of oracle bone script from distinct sources were collected and annotated. However, it is important to note that the annotation conventions for the oracle bone script may vary depending on the source. For instance, as shown in Table 2, some sources might use standard modern Chinese characters for annotation, while others might prefer Liding characters which are closer in form to the original oracle bone script. This leads to a scenario where images of oracle bone script that should belong to the same category are classified into different categories, creating redundant categories. Table 2 illustrates this with examples of duplicate annotations, where each row shows how the same oracle bone script image is categorized differently under the Modern Character Category and the Liding Character Category. To eliminate these redundancies, we integrate the data from different sources. For this purpose, we trained a widely-used unsupervised visual representation learning model MoCo [11], with oracle bone script images from all sources. Subsequently, all the oracle bone images were encoded into a feature vector by the model. As illustrated in Figure 6, by calculating the similarity of these feature vectors, we merged similar samples into the same categories. In this way, we were able to reduce the original 1,781 categories obtained from different sources to 1,588, eliminating repeated categories.
Validation
After undergoing all the procedures, we obtained a preliminary dataset. However, due to potential errors that might occur in the automated data acquisition and annotation process, we enlisted the expertise of OBS scholars to meticulously review and correct each sample. This thorough scrutiny culminated in the formation of the final HUST-OBC dataset.
Data Records
Source Type #Categories #Samples
Deciphered Undeciphered Deciphered Undeciphered
New Compilation of Oracle Bone Scripts Book 1,203 2,288 17,609 6,494
Oracle Bone Script: Six Digit Numerical Code Book 1,148 4,444 9,609 4,444
YinQiWenYuan Website 1,164 2,363 1,697 2,363
GuoXueDaShi Website 1,366 0 16,259 0
HWOBC Database 1,122 2,315 31,890 49,688
Total 1,588 9,411 77,064 62,989
Table 3: Statistic of oracle bone script data collected from various sources.
Figure 7: Histogram illustrating the distribution of sample counts across various categories in the HUST-OBC deciphered split.
The HUST-OBC comprises a total of 141,053 images sourced from five different origins, divided into deciphered and undeciphered sections. The deciphered section contains 77,064 images spanning 1,781 categories of individual characters, and the undeciphered section features 62,989 images across 9,411 categories of characters. Table 3 provides detailed statistics of the oracle bone script images obtained from these sources. Additionally, Figure 7 presents a distribution histogram showing the number of sample images per category in our dataset. It reveals that most categories have fewer than 10 sample images, with the largest category boasting over 300 images.
For efficient retrieval, the HUST-OBC is organized and stored by category names. Each image file is systematically named following the format <source>_<label>_<filename>, encapsulating its origin, category number, and sequence number, and is stored in folders named after their category number. For the deciphered categories, we have corresponding category numbers and the corresponding Chinese dictionary stored in a JSON file. Figure 8 demonstrates some deciphered and undeciphered oracle bone scripts from the HUST-OBC.
Figure 8: Samples of deciphered and undeciphered Oracle Bone Scripts from the HUST-OBC dataset.
Technical Validation
One of the primary objectives in creating the HUST-OBC is to facilitate future AI-assisted tasks in deciphering oracle bone scripts. To this end, we further assessed the quality of the dataset by employing it to train AI models. Specifically, we divided the deciphered section of the HUST-OBC dataset into a training set and a validation set in an 8:2 ratio, using them for training and testing in image classification tasks. The accuracy of image classification can reflect the quality of the dataset to some extent. If the images in the dataset are of poor quality or have many labeling errors, the classifier’s accuracy will be low, and vice versa. We employed the widely-used ResNet-50 [10] as the backbone network for training and ultimately achieved a classification accuracy of 94.3%, which validates the dataset’s quality and potential academic value. Table 4 shows the model’s recognition accuracy in some categories and provides example input images from different sources.
Usage Notes
The HUST-OBC is available as a compressed archive, comprising three distinct folders. These folders separately house images of oracle bone scripts. The first one is for those that have already been deciphered, the second one is for those still awaiting interpretation, and the third one is for unreliable data from GuoXueDaShi. Within each folder, subfolders are organized by categories, containing images of oracle bone scripts corresponding to their respective categories. For more information, please see (https://github.com/Pengjie-W/HUST-OBC).
Code Availability
OpenCV toolkit is used to detect the borders in scanned book pages, which is available at https://opencv.org/.
The models and code for Chinese OCR, MoCo, and ResNet50 for Validation is available at (https://github.com/Pengjie-W/HUST-OBC).
Sample Input from Different Sources
Ground-truth Category #Samples A B C D E Validation Accuracy
[Uncaptioned image] 307 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 83.61%
[Uncaptioned image] 295 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 96.61%
[Uncaptioned image] 209 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 100.00%
[Uncaptioned image] 199 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 97.50%
[Uncaptioned image] 190 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 97.37%
[Uncaptioned image] 180 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 97.22%
[Uncaptioned image] 179 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 100.00%
[Uncaptioned image] 106 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 95.24%
[Uncaptioned image] 93 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 100.00%
[Uncaptioned image] 63 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 100.00%
[Uncaptioned image] 55 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 100.00%
Table 4: Validation accuracy of ResNet-50 on partial categories of the HUST-OBC dataset.
Acknowledgements
The authors thank the Key Laboratory of Oracle Bone Script Information Processing, Ministry of Education, Anyang Normal University for providing ancient text data sources and review of the dataset construction.
Author contributions statement
Pengjie Wang conducted experiments on Automatic Annotation, Integration, and Technical Validation. Haisu Guan, Jinpeng Wan, Pengjie Wang, and Kaile Zhang collectively obtained data from books. Pengjie Wang and Kaile Zhang collectively crawled data from websites. Xinyu Wang analysed the results. Yuliang Liu provides guidance on the entire project. Xiang Bai provides laboratory resources. All authors reviewed the manuscript.
Competing interests
The corresponding author is responsible for providing a competing interests statement on behalf of all authors of the paper. This statement must be included in the submitted article file.